OpenStudy (anonymous):

Admission for a movie is $8 for children and $12 for adults. On a certain day, 3200 people attend the movie theater and $33,040 is collected. How many children and how many adults attended the movies

OpenStudy (anonymous):

8x+12y=33040 x+y= 3200 8x+8y= 25600 4y=7440 y=1860 3200- 1860=1340 1340 children, 1860 adults